Country A has 16,000 units of labor and can produce 2 goods, manufactures and food. A’s producers take 2 units of labor to produce one unit of manufactures and 8 units to produce one unit of food. Country B has 36,000 units of labor and takes 6 units of labor to produce one unit of manufactures and 5 units to produce one unit of food. What is the price of manufactures in terms of food at which A and B would respectively supply manufactures? What would A export in Adam Smith’s world? What is the amount it could supply?
Answer #1

Answer - Please find below mention answer.

A can make 1500 /5= 300 units of manufactures or 1500/6 = 250 units of food at the maximum B can make 1800/3=600 units of manufactures or 1800/6 = 300 food units Opportunity cost (OC) of food is 300/250=1.2manufactue in A and 600/300=2nits of manufactures in B OC of manufactures = 5/6 units of food in A and ½ in B So A has comparative advantage in food and will export food. B will supply manufactures. The world supply will depend on domestic demand for manufactures in B. If there is complete specialization, then B will supply 600 units in world market Manufactures 600 PPC of B ( Upper curve) 300 PPC of A 250 300 food.